# Characterizing the Efficiency of Distributed Training: A Power, Performance, and Thermal Perspective

### v1

@@ Acknowledgements.
This research was supported in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA, and by the AMD HPC Fund Research Cloud. We thank Eric Coulter and Michael Weiner for their assistance in provisioning access to this state of the art cluster. This work was partially supported by gifts from Google and AMD, and by ACE, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. We thank Matthieu Bloch for enabling this research. The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies or endorsements, either expressed or implied, of Georgia Tech.Appendix A Artifact Appendix

### v2

@@ Acknowledgements.
This research was supported in part through research cyberinfrastructure resources and services provided by the AI Makerspace and the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA, and by the AMD HPC Fund Research Cloud. We thank Eric Coulter and Michael Weiner for their assistance in provisioning access to the Georgia Tech AI Makerspace (of Technology, 2024; Forbes, 2025). This work was partially supported by gifts from Google and AMD, and by ACE, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. We thank Matthieu Bloch for enabling this research. The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies or endorsements, either expressed or implied, of Georgia Tech.Appendix A Artifact Appendix